# Revealing the Unseen: Guiding Personalized Diffusion Models to Expose Training Data

### v1

@@ 5.2 Generalization
(a) Ablation on $w^{\prime}$
(b) Ablation on $k$
Figure 5: Ablation Study on hype-parameters $w^{\prime}$ and $k$. We experiment on 4 classes of WikiArt dataset fine-tuning on SD (V1.4) with DreamBooth.
5.3 Ablation Study
In this section, we experiment with hyperparameters in Eq. 13, including the guidance scale $w^{\prime}$ and the correction term scale $k$. We experiment on 4 classes in WikiArt fine-tuning with DreamBooth. Results under LoRA are shown in Appendix Sec. D.
Figure 6: Visualization of generated images using FineXtract under different $w^{\prime}$ with one fixed starting noise $x_{T}$.
Guidance Scale $w^{\prime}$. The guidance scale $w^{\prime}$ is the most critical hyperparameter influencing extraction efficiency. If $w^{\prime}$ is too low, the guidance provided by fine-tuning methods is weakened. Conversely, if $w^{\prime}$ is too high, it often causes generation failures, resulting in unrealistic outputs (see visual examples in Fig. 6). As shown in Fig. 5, $w^{\prime}=3.0$ works well for both CFG and FineXtract when DMs are fine-tuned using DreamBooth.
Correction Term Scale $k$. In Eq. 13, we introduce a correction term $k\epsilon_{\theta}(x_{t},t)$ to address the inconsistency between CFG and model guidance. Although Eq. 13 indicates that $k$ should not be less than 0, our experiments suggest that setting $k=-0.02$ typically yields the best results, as illustrated in Fig. 5. This may be due to the complex interaction between $w^{\prime}$ and $k$, where $k$ can only be guaranteed to be greater than 0 if the true value of $w^{\prime}$ can be identified, which is often impractical in real-world scenarios.
@@ 5.4 FineXtract under Defense
The results presented in Tab. 5.4 illustrate how these methods can partially defend FineXtract, though at the cost of generation performance. Cutout and RandAugment indeed proves to be quite strong at defense. However, as shown in Fig. 7, the added transformations render the output images largely unusable, making them difficult to leverage in practice. Our results highlight that while these approaches may be partially effective in defense, there is a lack of research on how to fine-tune models on such transformed data while preserving the defensive effects. This remains an area for further investigation.
5.5 Real-world Results
Finally, we test our method on fine-tuned checkpoints available in the real world. We experiment on 10 checkpoints from HuggingFace where the corresponding training datasets are provided, allowing us to evaluate the effectiveness of our extraction method. Due to licensing restrictions, we only provide detailed information about the checkpoints with permissive licenses in Appendix Sec. 6. Quantitative results are shown in Table 4, where FineXtract consistently outperforms the baseline methods, increasing AS by at least 0.03 and doubling A-ESR in most cases.
@@ A.1 A Basic Scenario
$(\beta_{k}^{+}-\beta_{k}^{-})e_{i}^{T}=\sum_{j}\nabla L_{j}(\beta_{k}^{-}e_{0}^{T})\underbrace{e_{0}e_{i}^{T}}_{\text{a scalar}}.$ (16)
@@ A.3 Extension to More Optimizers and Approximation During Training with Multiple Linear Matrices
We can incorporate more information to reduce the effect brought by the error term. In practical scenarios, we have many linear projection layers that accept the same text embedding $e_{0}$ as input. For example, in SD, we may have about 20 such layers. Therefore, we can concatenate all of them together and try to find a decomposition considering them all. In other words, find a main vector for all matrices within the space $\mbox{span}\{e_{0,1},e_{0,2},\cdots,e_{0,77},\epsilon_{0,1},\epsilon_{0,2},\cdots,\epsilon_{0,77},\epsilon_{1,1},\epsilon_{1,2},\cdots,\epsilon_{0,77},\cdots,\epsilon_{K,1},\epsilon_{K,2},\cdots,\epsilon_{K,77}\}$, which corresponds to $\mbox{PCA}([(\beta_{0}^{+}-\beta_{0}^{-}),(\beta_{1}^{+}-\beta_{1}^{-}),\cdots,(\beta_{K}^{+}-\beta_{K}^{-})])$.
@@ A.4 Optimization
where $E(\cdot)$ is the frozen text feature embedding function. The challenge is that the prompt space is discrete, which turns the optimization problem into a hard-prompt finding task. To overcome this, we adopt the technique from Wen et al. (2024): during optimization, we project the current embedding onto a real prompt and use the gradient of this real prompt’s embedding to update our given embedding. Algorithm 1 illustrates this framework.
@@
  Input: Pre-trained Linear parameters $\beta^{-}$, fine-tuned Linear parameters $\beta^{+}$, input caption $c_{0}$, text encoder $E$, optimization iterations $N$, learning rate $\alpha$, projection function $E^{-1}$.
@@
  for $i=0$ to $N-1$ do
     $\hat{c}=E^{-1}(\hat{e})$
     Find gradient $\delta=\nabla_{E(\hat{c})}F(E(\hat{c}))$ based on this $\hat{c}$, $\beta^{+}$ and $\beta^{-}$ using Eq. 19
     Update $\hat{e}$rightarrow $\hat{e}+\alpha\delta$
@@ A.5 Experiment Result
Appendix B Computational Costs Analysis
Our method, as detailed in Sec. 4.1 and 4.2, employs guidance between two distinct models at each generation step without slowing down the generation speed compared to traditional CFG (which also requires guidance by forwarding the main UNet in DMs twice). However, GPU memory costs increase due to the need to load both the pretrained and fine-tuned models. In Tab. B, we present a demo experiment comparing the computational costs of FineXtract and CFG, using SD (v1.4) fine-tuned with DreamBooth and LoRA on the WikiArt dataset. The batch size is fixed at 5, and all experiments are conducted on a single A100 GPU.
[H] Generation GPU memory costs and time costs for per-image generation using FineXtract and CFG.
Metrics and Settings DreamBooth LoRA
Memory Costs (MB) Time Costs (s) Memory Costs (MB) Time Costs (s)
FineXtract 15812 3.4 15894 3.6
CFG 12096 3.4 12226 3.6
Appendix C Results with No Prompts Provided
@@
Appendix D Abaltion Study of $w^{\prime}$ under LoRA Fine-tuning
@@
Figure 9: Ablation Study on hype-parameters $w^{\prime}$.
Appendix E Finetuning Details
@@
Appendix F Real-World Experiments Setup

### v3

@@ 5.4 FineXtract under Defense
The results presented in Tab. 4 illustrate how these methods can partially defend FineXtract, though at the cost of generation performance. Cutout and RandAugment indeed proves to be quite strong at defense. However, as shown in Fig. 7, the added transformations render the output images largely unusable, making them difficult to leverage in practice. Quantitative measurements of this unusability are provided in Appendix Section I. Our results highlight that while these approaches may be partially effective in defense, there is a lack of research on how to fine-tune models on such transformed data while preserving the defensive effects. This remains an area for further investigation.
@@
5.5 Real-world Results
Finally, we test our method on fine-tuned checkpoints available in the real world. We experiment on 10 checkpoints from HuggingFace where the corresponding training datasets are provided, allowing us to evaluate the effectiveness of our extraction method. Due to licensing restrictions, we only provide detailed information about the checkpoints with permissive licenses in Appendix Sec. 8. Quantitative results are shown in Tab. 5, where FineXtract consistently outperforms the baseline methods, increasing AS by at least 0.03 and doubling A-ESR in most cases.
@@ 6 Conclusion
Acknowledgments
We sincerely thank Peishen Yan, Yuxiang Lu, Xirui Li, and Suizhi Huang for their valuable feedback and suggestions on this paper.
Impact Statement
This paper explores methods for extracting fine-tuned data from publicly available checkpoints. The proposed approach offers a dual impact: it can serve as a legal tool to detect unauthorized data usage, while also raising awareness about the risks of data leakage for those who share fine-tuned models online. These insights are intended to contribute to advancing responsible practices in model sharing and checkpoint distribution.
From an ethical perspective, this work highlights the need for caution and transparency in model fine-tuning and sharing, particularly concerning sensitive or proprietary data. The societal consequences of this work include promoting accountability in data usage while underscoring the importance of copyright protection and privacy in machine learning. However, we acknowledge that potential attackers may misuse the techniques described in this paper, which could pose significant threats. By revealing these vulnerabilities ahead of time, this paper aims to inspire the development of future defense methods to safeguard against such risks.
@@ A.1 A Basic Scenario
$(\beta_{k}^{+}-\beta_{k}^{-})e_{i}^{T}=\sum_{j}\nabla L_{j}(\beta_{k}^{-}e_{0}^{T})\underbrace{e_{0}e_{i}^{T}}_{\text{\scriptsize a scalar}}.$ (16)
@@ A.3 Extension to More Optimizers and Approximation During Training with Multiple Linear Matrices
We can incorporate more information to reduce the effect brought by the error term. In practical scenarios, we have many linear projection layers that accept the same text embedding $e_{0}$ as input. For example, in SD, we may have about 20 such layers. Therefore, we can concatenate all of them together and try to find a decomposition considering them all. In other words, find a main vector for all matrices within the space $\mbox{span}\{e_{0,1},e_{0,2},\cdots,e_{0,77},\epsilon_{0,1},\epsilon_{0,2},\cdots,\epsilon_{0,77},\epsilon_{1,1},\epsilon_{1,2},\cdots,\epsilon_{1,77},\cdots,\epsilon_{K,1},\epsilon_{K,2},\cdots,\epsilon_{K,77}\}$, which corresponds to $\mbox{PCA}([(\beta_{0}^{+}-\beta_{0}^{-}),(\beta_{1}^{+}-\beta_{1}^{-}),\cdots,(\beta_{K}^{+}-\beta_{K}^{-})])$.
@@ A.4 Optimization
where $E(\cdot)$ is the frozen text feature embedding function. The challenge is that the prompt space is discrete, which turns the optimization problem into a hard-prompt finding task. To overcome this, we adopt the technique from Wen et al. (2024a): during optimization, we project the current embedding onto a real prompt and use the gradient of this real prompt’s embedding to update our given embedding. Algorithm 1 illustrates this framework.
@@
Input: Pre-trained Linear parameters $\beta^{-}$, fine-tuned Linear parameters $\beta^{+}$, input caption $c_{0}$, text encoder $E$, optimization iterations NN, learning rate $\alpha$, projection function $E^{-1}$.
@@
for $i=0$ to N1N-1 do
[…]
  Find gradient $)\delta=\nabla_{E(\hat{c})}F(E(\hat{c}$)) based on this $\hat{c}$, $\beta^{+}$ and $\beta^{-}$ using Eq. 19
  Update $\hat{e} \rightarrow$ $\hat{e}+\alpha\delta$
@@ A.5 Experiment Result
Experiment results under longer training prompt are discussed in Sec. J.
Table 7: Extraction comparison on fine-tuned SD v1.4 checkpoints using DreamBooth under different caption settings.
Extraction Method AS $\uparrow$ A-ESR0.6 $\uparrow$
FineXtract (Full Caption) 0.501 0.35
FineXtract (Extracted Caption) 0.314 0.15
FineXtract (Empty Caption) 0.192 0.00
CFG + Clustering (Full Caption) 0.434 0.23
CFG + Clustering (Extracted Caption) 0.308 0.08
Direct Text2Img + Clustering (Empty Caption) 0.146 0.00
We further evaluate our method using extracted captions (3 words in length) as shown in Tab. 7. While the success rate decreases compared to using full captions, it remains substantially higher than extraction without any caption information, confirming the effectiveness of our caption extraction approach. Moreover, FineXtract consistently outperforms the baseline under the extracted caption setting.
@@ Appendix B Results with No Prompts Provided
Appendix C Ablation Study of $w^{\prime}$ under LoRA Fine-tuning
@@
Figure 9: Ablation Study on hyper-parameters $w^{\prime}$.
Appendix D Finetuning Details
@@
Appendix E Real-World Experiments Setup
@@
Appendix F Experiment on Mixture of Dataset
F.1 Mixture of DreamBooth and WikiArt Datasets
We constructed a new dataset with 10 classes, each containing 5 images from DreamBooth and 5 from WikiArt. As reported in Tab. 9, this mixture led to mutual decreases in the fine-tuned model’s fidelity (measured by DINO score (Ruiz et al., 2023)), image quality (measured by CLIP-IQA (Wang et al., 2023)), and extraction rate. Therefore, extracting training images from this model becomes more challenging. Nonetheless, FineXtract still significantly outperforms the baseline despite the fact that its performance partially drops compared to the original cases, further verifying its generality and effectiveness.
Table 9: Comparison of the performance of FineXtract and the baseline in fine-tuning data extraction under separated and mixed data scenarios using DreamBooth and WikiArt datasets. We observe mixed data makes extraction more challenging while also making it harder for diffusion models to learn the given data, as evidenced by lower fidelity (DINO) and reduced image quality (CLIP-IQA).
[…]
Separated Data CFG+Clustering 0.525 0.45 0.533 0.697
FineXtract 0.572 0.55
Mixed Data CFG+Clustering 0.457 0.08 0.266 0.447
FineXtract 0.480 0.18
F.2 Mixture of Different Styles in WikiArt Datasets
We conducted experiments with an increasing number of classes and found that while it reduces our method’s performance, we still significantly outperform the baseline. As shown in Tab. 10, we observe that as the number of styles increases, the extraction success rate decreases. Additionally, we evaluated the model’s ability to learn the input distribution and found that a higher number of classes leads to lower fidelity (DINO) and image quality (Clip-IQA), reflecting the model’s difficulty in learning the fine-tuning data distribution accurately.
Table 10: Comparison of the performance FineXtract and baseline in fine-tuning data extraction under dataset mixed from different number of classes of WikiArt dataset. Mixed data with different classes makes model learning poorer, causing lower fidelity (DINO) and image quality (Clip-IQA), which in turn makes extraction more challenging.
[…]
1 Artist Per Class CFG+Clustering 0.396 0.11 0.458 0.525
FineXtract 0.449 0.22
2 Artist Per Class CFG+Clustering 0.390 0.20 0.387 0.441
FineXtract 0.436 0.20
5 Artist Per Class CFG+Clustering 0.353 0.15 0.343 0.478
FineXtract 0.388 0.23
Appendix G Ablation Study on Model Guidance
Figure 10: Ablation study on different $w^{\prime}$ dealing with model guidance, CFG, and FineXtract. We experiment on 4 classes of WikiArt dataset.
We provide further discussion by comparing model guidance using conditional DM, CFG, and a combination of CFG and model guidance (with $k=0$). As shown in Fig. 10, model guidance using conditional DM achieves close performance to the combination method and significantly outperforms CFG alone, suggesting model guidance’s dominance and a potential misalignment with CFG, affecting the parameter $k$.
Appendix H Experiment Results on FLUX.
We update the results using FLUX.1 [dev]  [^8] https://huggingface.co/black-forest-labs/FLUX.1-dev. The DreamBooth fine-tuning on FLUX.1 [dev] requires more than 100GB per GPU, which we are currently unable to run. However, we have conducted experiments in the LoRA scenario using the official scripts provided by Diffusers. The training iterations are fixed at 150, as suggested by the repository, and other hyperparameters remain the same as those in the repo. As shown in Tab. 11, we compare our method with the baseline (CFG), and both methods show the best performance when the guidance strength is set to 3.0, with the improvement being consistent.
Table 11: Extraction comparison between the baseline and our method on fine-tuned FLUX checkpoints using LoRA.
[…]
CFG+Clustering 0.568 0.458 0.525
FineXtract 0.507 0.460 0.487
Appendix I More Assessments for Pre-processing Defense
We attempted to quantify usability using a no-reference image quality score and the image fidelity measurements. We follow previous works (Ruiz et al., 2023), using DINO for image fidelity measurements. For image quality measurements, we use Clip-IQA (Wang et al., 2023). As shown in Tab. 12, we found that the CutOut and RandAugment degraded Clip-IQA by 6.1% and 4.6% in the original dataset, and generated images experienced around 3.8% and 4.6% degradation. These extent are close in the sense that the degration brought by the transformation largely preserves in the generation process. For RandAugment, the image fidelity also largely degrades. This suggests that the preprocessing steps applied to the images (removing a square section or adding high contrast to the input images) largely persist in the generated output images, hindering extraction methods from obtaining high-quality images while sacrificing generation quality of diffusion models themselves.
Table 12: Image fidelity and quality for checkpoints under different defenses. DINO measures image fidelity with respect to the training dataset, and thus it is not applicable when dealing with source images. We use “N.A.” to denote this scenario.
Source Images Generated Images
Defenses $uparrow DINO\uparrow Clip-IQA\uparrow$
No Defense N.A. 0.568 0.458 0.525
CutOut N.A. 0.507 0.460 0.487
RandAugment N.A. 0.522 0.435 0.479
In summary, there exists a trade-off between image quality and defensive effects for these preprocessing methods. We leave the accurate modeling of this trade-off and further improvement as an interesting future work.
Appendix J Experiment on Extracting Prompts under Different Prompt Length $W$
We further experiment on different scenarios involving fine-tuning with longer training prompts. Specifically, we use GPT-4 to expand the original prompts into longer versions by adding information related to the author while avoiding overly specific details to prevent mismatches with the input image. Tab. 13 shows our extraction results. With longer text, extracting detailed information becomes more challenging. However, our algorithm can still identify specific details, such as the artist’s name, even with extended text. When the text length continue to increase, certain information becomes harder to extract, though such cases are rare in few-shot fine-tuning.
Table 13: Experimental results of our caption extraction algorithm using the L2-PGD attack with 1000 iterations starting from a random prompt. This extraction costs approximately 1.5 to 2 minutes per sample on a single A100 GPU. The results demonstrate that key information, such as the artist’s name, can largely be inferred when the training caption is not too long. Correctly inferred key-word parts are shown in bold. Related information extracted is shown in italic.
Extended Prompt Training Caption Extracted Caption ($\leq$ 7 words) An insight into the Post-Impressionist style with expressive, emotional brushwork and rich colors, inspired by the unique techniques of Vincent van Gogh. vangogh impressionism class impressionist paintings acqu Exploring the art style of Fauvism, characterized by vivid, bold colors and dynamic brushstrokes, as seen in the works of Henri Matisse. henri matisse whose explores whose colourful stures An exploration of the High Renaissance style, marked by balance, realism, and anatomical accuracy, capturing the essence of Leonardo da Vinci’s art. cws inaccurate anatomy leonardo deus aron onda A look at Impressionism, noted for soft, light-filled scenes and gentle brushstrokes, inspired by the atmospheric beauty in Claude Monet’s paintings. impressionism wgleagues monet het commence
Excessively Extended Prompt Training Caption Extracted Caption ($\leq$ 7 words) A deeper look into Post-Impressionism, emphasizing Vincent van Gogh’s expressive, emotional brushwork and contrasting colors. His techniques convey a personal, intense view of the world, moving beyond the lighter tones of Impressionism. post impressionism vincent gogh pcdimonet opposite vangogh In-depth exploration of Fauvism’s vivid colors and bold, expressive brushstrokes that bring emotions to life, heavily inspired by the renowned techniques of Henri Matisse. His approach captures the vibrant essence of Fauvism, using striking colors to convey feeling. congratulations adrian edgar forum awaits muromain destination An insightful exploration into the High Renaissance, highlighting the balance, anatomical realism, and idealism of Leonardo da Vinci’s creations. This style embodies the Renaissance’s focus on perfection, with meticulously detailed compositions and harmonious proportions. retains vehicles marriott alcatraspberries tuesdaythoughts shua A detailed depiction of Impressionism, capturing transient beauty with soft brushstrokes and delicate light, influenced by Claude Monet’s signature style. His work portrays nature in harmonious, gentle colors that evoke calmness. impressionism excellent depiction effectively arranged unexrepresented explained